OpenStudy (anonymous):

simplify completely (x^2-3x-54/x^2-18x+81)*(x^2+12x+36/x+6) please :) I'm a little stuck

OpenStudy (unklerhaukus):

\[ \frac{x^2-3x-54}{x^2-18x+81}\times\frac{x^2+12x+36}{x+6}\]

OpenStudy (anonymous):

yes! that is much clearer than the way I wrote it, but can you explain what steps you use ti solve it?

OpenStudy (anonymous):

i mean simplify***

OpenStudy (unklerhaukus):

ok so there are two numerators (the bit a the top of a fraction) and two denominators (the bit at the bottom of the fraction), factorize each of these in to something like (x+...)(x+...) {the x+6 can not be factorized further then look to cancel common bracketed terms that are in the numerator and the denominator

OpenStudy (anonymous):

so the final outcome of (x+6)^2 / x-9 would make sense right ?

OpenStudy (unklerhaukus):

that looks great

OpenStudy (anonymous):

thanks so much!